I have six more dimes then quarters in my pocket. the total amount of money I have in my pocket is $2.00. how many quarters do i have? how many dimes do i have?
Answer by mananth(16949)

let quarters be x
dimes = x+6
..
total money = 200 cents
25x+10(x+6)=200
25x+10x+60=200
35x+60=200
-60
35x+60-60=200-60
35x = 140
/35
x=140/35
x=4 quarters
dimes = x+6 =10